### v1

# Kinetic theory of decentralized learning for smart active matter

Abstract
Smart active matter describes agents which can process information according to their individual policy to solve predefined tasks. We introduce a theoretical framework to study a decentralized learning process in which agents can locally exchange their policy to optimize their reward and thus attain target macroscopic states. We use our formalism to derive explicit hydrodynamic equations for the policy dynamics. We apply the theory to two different microscopic models where policies correspond either to fixed parameters similar to evolutionary dynamics, or to state-dependent controllers known from the field of robotics. We find good agreement between theoretical predictions and agent-based simulations. By deriving fundamental control parameters and uncertainty relations, our work lays the foundations for a statistical physics analysis of decentralized learning.
Decentralized optimization processes have been intensively studied in interdisciplinary fields using methods inspired from statistical physics, including the study of evolutionary and population dynamics in biology [1, 2, 3, 4] and the investigation of social processes [5], such as opinion dynamics [6]. In addition, decentralized adaptation raises a growing interest in robotics and engineering [7, 8, 9, 10, 11], in particular to find more stable policies than typical centralized protocols [9, 10]. Robustness and flexibility may be enhanced by drawing inspiration from collective animal behavior [12]. Although standard experiments on robotic swarms are in practice often limited to less than a hundred robots, numerical simulations may account for much larger swarms, calling for a statistical physics approach to describe assemblies of small robots interacting through information exchange [10].
The idea of experimentally realizing large swarms of microrobots [13, 14, 15] as a way to build responsive or programmable metamaterials [16, 17, 18, 19] is also at the core of the emerging field of smart active matter [20, 21, 22, 23, 24]. Along this soft robotics perspective, one tries to build large numbers of extremely simplified microrobots [25, 26, 27, 28], whose size ranges from the granular scale [29, 30] down to the colloidal one [31, 19], and whose main features result from the physical properties of the soft material they are made of [32, 33, 34, 35]. These soft microrobots may integrate simple computation capabilities [36], or may be controlled by an external computer through machine learning using, e.g., reinforcement learning techniques [37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. Although their starting points differ, both hard and soft robotics aim at building large assemblies of autonomous microrobots with communication and computation capabilities leading to adaptative collective behaviors [10, 11, 22, 23, 26, 47, 48]. This goal is also shared by the emerging topic of biological metamaterials, consisting for instance of insects aggregates [49, 50].
The ongoing experimental and numerical development of smart active matter calls for appropriate theoretical approaches to describe its collective properties. Early contributions in this direction include models of individual [51, 52, 53, 54, 55] or collective navigation of active particles [56, 57], or of pedestrians in a dense crowd [58, 59]. Stochastic thermodynamic may also be relevant to describe smart active matter [60, 61]. Yet, apart from some first steps [60], a theoretical framework to describe the collective behavior of adaptive agents exchanging information with their neighbors, as found in experiments with small robots [7, 8, 9, 10, 11], is still lacking.
In this Letter, we develop a statistical physics framework based on kinetic theory to describe decentralized learning in smart active matter with local information exchange. In contrast to usual active matter [62, 63, 64], smart active matter is made of agents having the ability to communicate, to respond to external stimuli and to adapt their behavior to attain targeted collective properties. The behavior of individual agents is determined by a ‘policy’, i.e., a set of rules controlling their dynamics, whose parameters can be tuned. Among the modeling approaches used to characterize active matter, kinetic theories have played an important role by providing a framework to derive macroscopic hydrodynamic equations from microscopic collision rules [65, 66, 67, 68, 69]. Here, we describe decentralized learning via the exchange of policies between neighboring agents, akin to collision rules. Such binary exchange rules set the stage for a kinetic theory framework, within which we systematically derive macroscopic hydrodynamic equations for the time-evolution of policies starting from agent-based models.
Figure 1: Sketch of the decentralized learning procedure described in the manuscript.
Agent-based model of decentralized learning – We consider a generic system of $N$ agents in $d$ dimensions representing, e.g., small robots [10]. Agents are located at position $\bm{x}$, move at velocity $\bm{v}$ with an overdamped dynamics, and diffuse with a positional diffusion coefficient $D_{0}$. Agents sense their environment (e.g., external fields or properties of neighboring particles) to control their dynamical rules. The set of adjustable parameters controlling the dynamics, which are specific to each agent, is here generically called the agent’s policy $\mathcal{P}$. In the examples below, the policy either corresponds to the agent’s angular diffusion coefficient, or to the response of the self-propulsion speed to an external stimulus. The state $\mathcal{S}$ of an agent gathers internal physical variables (e.g., orientation, self-propulsion velocity, etc.) and sensors outputs such as measured light intensity. Agents can then process observed states and extract the signal $G(\mathcal{S})$, which fluctuates with a characteristic frequency $\tau_{G}^{-1}$. The signal can be stored in the ‘memory’ $\mathcal{M}$ via the running average $\mathcal{M}(t)$ through
$\frac{\text{d}\mathcal{M}}{\text{d}t}=-\frac{1}{\tau_{\mathcal{M}}}\,\big{(}\mathcal{M}-G(\mathcal{S})\big{)},$ (1)
on the memory time scale $\tau_{\mathcal{M}}$. All quantities $\mathcal{S}$, $\mathcal{M}$, and $\mathcal{P}$ may be multi-component vectors. To lighten the presentation of the general formalism, we mostly stick to the scalar case, but the multi-component vector formalism is presented in the supplemental material (SM) [70].
To enable learning, agents can exchange information about the efficiency of their individual policies to reach a targeted collective behavior. This efficiency is quantified by a reward function $\mathcal{R}(\mathcal{M})$, which rewards agents following the targeted collective behavior. Comparing its own reward with other agents enables each agent to learn improved policies (see Fig. 1 for an illustration). The overall goal of each agent is to learn and optimize a policy $\mathcal{P}$ maximizing the reward function. Agents are therefore randomly selected with a rate $\lambda_{E}=\tau_{E}^{-1}$, called the exchange rate ($\tau_{E}$ is the exchange time). Once selected, an agent $i$ iterates over all neighbors $n$, chosen in random order within the interactions radius $r_{c}$. For each selected neighbor $n$, agent $i$ becomes a teacher (and agent $n$ the student) with a probability $p_{T}(\mathcal{R}_{i},\mathcal{R}_{n})$ given by
$p_{T}(\mathcal{R}_{i},\mathcal{R}_{n})=\frac{1}{2}\left(1+\tanh\big{[}\alpha_{T}\big{(}\mathcal{R}_{i}-\mathcal{R}_{n}\big{)}\big{]}\right),$ (2)
where $\mathcal{R}_{i}=\mathcal{R}(\mathcal{M}_{i})$ and $\mathcal{R}_{n}=\mathcal{R}(\mathcal{M}_{n})$, and $\alpha_{T}$ is a parameter. Otherwise, agent $n$ is the teacher and agent $i$ is the student. In either case, the student updates its policy $\mathcal{P}_{S}$ and memory $\mathcal{M}_{S}$ by taking the corresponding values of the teacher, $\mathcal{P}_{S}:=\mathcal{P}_{T}$ and $\mathcal{M}_{S}:=\mathcal{M}_{T}$. The memory of the student is updated together with the policy to make its reward consistent with the updated policy [10]. The parameter $\alpha_{T}$ in Eq. (2) controls the sharpness of $p_{T}$, and may be interpreted as the precision with which agents evaluate and communicate the reward $\mathcal{R}$. To enhance adaptability, we also include a small rate of spontaneous policy changes. By analogy with evolutionary processes in biology, we call these spontaneous changes ‘mutations’ [71] and model them as a diffusion process in policy space with diffusion coefficient $D_{M}$.
Kinetic theory of decentralized learning – To obtain a statistical description of the above agent-based model and derive the time-evolution for the statistics of policies $\mathcal{P}$, we introduce a kinetic theory framework. Learning events between two agents are treated by analogy with binary collisions in usual kinetic theory. In this framework, an assembly of a large number $N$ of agents is described by a single-agent phase-space density $f(\mathcal{S},\mathcal{M},\mathcal{P},\bm{x},t)$ characterizing the probability to find an agent at time $t$ at position $\bm{x}$, in state $\mathcal{S}$, with memory $\mathcal{M}$ and policy $\mathcal{P}$. The time-evolution of $f(\mathcal{S},\mathcal{M},\mathcal{P},\bm{x},t)$ formally reads
$\frac{\partial f}{\partial t}=\mathcal{I}_{\rm phys}[f]+\mathcal{I}_{\rm sens}[f]+\mathcal{I}_{\rm mem}[f]+\mathcal{I}_{\rm pol}[f].$ (3)
The term $\mathcal{I}_{\rm phys}[f]$ describes the physical part of agents dynamics, e.g., their motion, angular dynamics and possibly physical interactions, and $\mathcal{I}_{\rm sens}[f]$ defines the dynamics of the sensor output. The term $\mathcal{I}_{\rm mem}[f]$ in Eq. (3) encodes the memory dynamics given in Eq. (1). Finally, the term $\mathcal{I}_{\rm pol}[f]$ accounts for the dynamics of the policy $\mathcal{P}$, which includes diffusive mutations and communication-driven decentralized learning,
$\mathcal{I}_{\rm pol}[f]=D_{M}\frac{\partial^{2}f}{\partial\mathcal{P}^{2}}+\mathcal{I}_{\rm learn}[f].$ (4)
The learning term $\mathcal{I}_{\rm learn}[f]$ represents the exchange of information between neighboring agents regarding their policies and the corresponding rewards. Mathematically, the term $\mathcal{I}_{\rm learn}[f]$ is a bilinear integral term accounting for collision-like, binary learning events. The explicit, but lengthy, expression of $\mathcal{I}_{\rm learn}[f]$ is reported in the End Matter (EM) section. We assume in this manuscript a time scale separation $\tau_{G}\ll\tau_{\mathcal{M}}\ll\tau_{E}$, i.e., a hierarchy of time scales separating signal dynamics $\tau_{G}$, memory $\tau_{\mathcal{M}}$ and exchange rate $\tau_{E}$. As a result, information advection by particle motion dominates over information propagation due to the finite communication range between agents. Within this approximation, our theory is thus purely local, however, it can be generalized to take into account the non-locality of communications or overlapping time scales.
To characterize the policy statistics, we assume that the reduced phase-space density
[…]
is approximately Gaussian and can be described by its lowest-order moments. We also define the policy-dependent average memory,
[…]
The reduction steps from $f(\mathcal{S},\mathcal{M},\mathcal{P})$ to $\varphi_{0}(\mathcal{P})$ and $\mu_{\mathcal{M}}(\mathcal{P})$ are described in the EM section. To keep the problem tractable, we expand $\mu_{\mathcal{M}}(\mathcal{P},\bm{x},t)$ around a predefined policy value $\mathcal{P}_{E}$, which should be a rough estimate of the optimal policy, yielding (dropping $\bm{x}$ and $t$)
$\mu_{\mathcal{M}}(\mathcal{P})=\mu_{\mathcal{M}}^{(0)}+\mu_{\mathcal{M}}^{(1)}(\mathcal{P}-\mathcal{P}_{E})+\dots$ (7)
[…]
For spatially uniform fields $\rho(t)$, $\mu(t)$ and $\sigma^{2}(t)$, the evolution equations take the form,
$\frac{d\mu}{dt}$ $=\rho\sigma^{2}F_{1}\left(\mu,\sigma^{2};\mu_{\mathcal{M}}^{(n)}\right),$ (10)
$\frac{d\sigma^{2}}{dt}$ $=2D_{M}-\rho\sigma^{4}F_{2}\left(\mu,\sigma^{2};\mu_{\mathcal{M}}^{(n)}\right),$ (11)
where $F_{m}\big{(}\mu,\sigma^{2};\mu_{\mathcal{M}}^{(n)}\big{)}$ are model-specific polynomials [70] depending on the parameters $\mu_{\mathcal{M}}^{(n)}(t)$ introduced in Eq. (7), which can be determined independently. The closed equations (10), (11) for the policy dynamics are one of the major results of this work. The full hydrodynamic equations including space-dependent fields are derived for two specific models in the SM [70]. Importantly, we can already draw some non-trivial general conclusions from Eqs. (10) and (11): (i) The derivative $\dot{\mu}(t)$ is proportional to the density $\rho(t)$ highlighting the many-body character of policy dynamics. Similarly, it is proportional to the diversity $\sigma^{2}(t)$, thus vanishing density or diversity imply vanishing adaptation rate $\dot{\mu}(t)$, consistent with Fisher’s fundamental theorem of natural selection [72]. (ii) On time scales $t>\tau_{\mathcal{M}}$ with convex fitness function, we have $F_{2}(\mu(t),\sigma^{2}(t);\mu_{\mathcal{M}}^{(n)}(t))>0$ and thus $\sigma^{2}(t)$ decays monotonically in the absence of mutations. (iii) This emphasizes the importance of mutations $D_{M}$ to maintain a sufficient level of diversity and thus enable efficient optimization. Qualitatively, mutations therefore play a role similar to driving in athermal physical systems, as they break detailed balance and maintain a steady-state dynamics at large time.
Figure 2: Adaptation of the microswimmer model to a constant target velocity $V_{T}=0.85$: Results of kinetic theory (KT) and agent-based simulations (AS) without mutations ($D_{M}=0$) and with mutations ($D_{M}=0.1$). The vertical arrow indicates $\tau_{\mathcal{M}}=1$. (a) Average policy (rotation diffusion) $\mu(t)$; (b) diversity $\sigma^{2}(t)$. The black dashed-dotted curve is Eq. (12) fitted to the AS data with $D_{M}=0.1$.
Phototactic, diffusive agents – As a first application of the above framework, we study a model of phototactic microswimmers ($d=2$) with decentralized learning capabilities. Each swimmer moves at speed $v_{0}$ in a direction given by its orientation angle $\theta$ in the $(x,y)$-plane, which defines the state $\mathcal{S}=\theta$. As for active Brownian particles [73], orientation can change randomly via rotational diffusion with strength $D_{\theta},$ but also via tumbling events towards the $x-$direction (direction of the light source) [74]. Physical interactions between microswimmers are neglected. The target state is a macroscopic flow along the $x$-axis in which the ensemble-averaged velocity $\bar{V}_{x}(D_{\theta})=\langle V_{x}\rangle_{D_{\theta}}$ is equal to $V_{T}<v_{0}$. This target state is achieved via a decentralized learning procedure in which each agent measures its velocity $G(\mathcal{S})=V_{x}=v_{0}\cos\theta$ and adapts its rotational diffusion coefficient $\mathcal{P}=D_{\theta}$ (i.e., its policy) by maximizing the reward function, $\mathcal{R}(\mathcal{M})=1-(\mathcal{M}-V_{T})^{2}$. Importantly, although physical interactions are neglected, learning leads to collective effects. The asymptotic solutions of Eqs. (10) and (11) (detailed derivations in [70]) read,
$\sigma^{2}(t)$ $=\sqrt{\frac{2D_{M}}{\lambda_{0}}}\frac{1}{\tanh(\sqrt{2D_{M}\lambda_{0}}(\tau_{0}+t))}$ (12)
$\mu(t)$ $=D_{\theta,T}-\frac{\mu_{0}\sqrt{2D_{M}\lambda_{0}}}{\sinh(\sqrt{2D_{M}\lambda_{0}}(\tau_{0}+t))},$ (13)
which boil down to $\sigma^{2}(t)=[\lambda_{0}(\tau_{0}+t)]^{-1}$ and $\mu(t)=D_{\theta,T}-\mu_{0}/(\tau_{0}+t)$ for $D_{M}\to 0$. The parameter $\lambda_{0}=4\lambda_{E}\alpha_{T}\rho{V_{E}^{\prime}}^{2}$ is a characteristic rate, where $V_{E}^{\prime}={\partial\bar{V}_{x}(D_{\theta})}/{\partial D_{\theta}}|_{D_{\theta}=D_{\theta,E}}$ is the response of the average state to policy changes. The parameters $\tau_{0}$ and $\mu_{0}$ depend on initial conditions, and $D_{\theta,E}=\mathcal{P}_{E}$ is the predefined policy used in Eq. (7). When $D_{M}>0$, $\sigma^{2}(t)$ converges to a plateau value $\sigma^{2}_{\infty}=\sqrt{2D_{M}/\lambda_{0}}$ when $t\to\infty$, while $\mu(t)$ converges to the optimal policy $D_{\theta,T}$. The convergence of both $\sigma^{2}(t)$ and $\mu(t)$ occurs over the learning time $\tau_{L}=(2D_{M}\lambda_{0})^{-1/2}$. Importantly, our theoretical analysis reveals two important control parameters: the rate $\lambda_{0}$ and the mutation strength $D_{M}$. The effect of the mutation rate $D_{M}$ is characterized by the ‘uncertainty relation’ $\tau_{L}\sigma^{2}_{\infty}=\lambda_{0}^{-1}$, where both $\tau_{L}$ and $\sigma^{2}_{\infty}$ depend on $D_{M}$, while $\lambda_{0}$ is independent of $D_{M}$. The uncertainty relation states that quick convergence and adaptability (i.e., a small $\tau_{L}$) obtained by increasing the mutation rate $D_{M}$ comes at the cost of having large fluctuations around the target policy (i.e., a large $\sigma^{2}_{\infty}$), in turn implying collective fluctuations around the target state, since $\mathrm{Var}(N^{-1}\sum_{i=1}^{N}V_{x,i})\propto\sigma^{2}_{\infty}/N$. Predictions of the kinetic theory are successfully compared to agent-based simulations in Fig. 2. Equation (12) can be used to fit the long-time behavior of $\sigma^{2}(t)$ on agent-based simulations data (see Fig. 2) to determine $\lambda_{0}$ and $D_{M}$, as on experimental data where such parameters are unknown. We get $D_{M}^{\rm fit}=0.094\pm 0.002$ and $\lambda_{0}^{\rm fit}=(2.65\pm 0.01)\cdot 10^{-4}$, which compare well to the agent-based model values $D_{M}=0.1$ and $\lambda_{0}=2.655\cdot 10^{-4}$.
Figure 3: Adaptation of the microswimmer model to a space-dependent target $V_{T}(x)=0.85+0.05\sin(2\pi x/L)$ (black-dashed line). Results obtained by kinetic theory (KT) and agent-based simulations (AS) are shown for different mutation rates $D_{M}.$
We now consider a space-dependent target velocity $V_{T}(x)=V_{T}^{s}+V^{\Delta}_{T}\sin(2\pi x/L)$ in a system of size $L$ with periodic boundary conditions. The reward function thus also becomes space-dependent, $\mathcal{R}(\mathcal{M},x)=1-\big{(}\mathcal{M}-V_{T}(x)\big{)}^{2}$. At leading order, the steady-state average policy $\mu(x)$ satisfies [70],
$-V_{T}^{s}\nabla\mu(x)-\sqrt{2D_{M}\lambda_{L}}(\mu(x)-D_{\theta,T}(x))=0,$ (14)
where $D_{\theta,T}(x)$ is the optimal policy maximizing the local reward. We find for the average local velocity
$\langle V_{x}(x)\rangle=V_{T}^{s}+V^{\Delta}_{x}(\phi)\sin(2\pi x/L-\phi),$ (15)
with a phase $\phi$ given by $\tan(\phi)=2\pi V_{T}^{s}/(\sqrt{2D_{M}\lambda_{0}}L)$ and an amplitude $V^{\Delta}_{x}(\phi)=V^{\Delta}_{T}\cos(\phi)$. Agents with small mutation rates travel across the system faster than they can adapt to the local target $V_{T}(x)$, therefore adopting an average speed $V_{T}^{s}.$ In Fig. 3, we compare this predicted behavior by kinetic theory to agent-based simulations and find very good agreement for both the mutation-rate dependent amplitude $V^{\Delta}_{x}(D_{M})$ and the phase $\phi(D_{M})$. Our results emphasize the importance of the mutation rate $D_{M}$ to maintain diversity and therefore being able to continuously adapt to space- and time-dependent targets.
Figure 4: Adaptation of the light-sensing robot model to maximize collected light: Kinetic theory (KT) and agent-based simulations (AS) without mutations ($D_{M}=0$) and with mutations ($D_{M}=10^{-3}$). (a) Average policy $\mu(t)=\langle\chi\rangle$; (b) Average reward ¯$\bar{\mathcal{R}}(\chi)$ for robots with a given policy $\chi$.
Light-sensing robots – We also study a model featuring robots ($d=1$) which attempt to maximize light collected from a heterogeneous light field $I(x)$. The state of a robot is $\mathcal{S}=(I,v)$, where $I$ is the local light intensity measured by a sensor, and $v$ is the robot velocity (diffusion is neglected, $D_{0}=0$). The memory $\mathcal{M}$ averages the measured light intensity according to Eq. (1), with $G(\mathcal{S})=I$. A state-dependent controller adapts the robot’s speed $v(I)=v_{0}-\chi I$, keeping $v(I)>v_{\rm min}$. The parameter $\chi$ thus defines the sensitivity with which the robot reacts to the external light. Robots adapt the policy $\mathcal{P}=\chi$ to maximize the collected light, using a reward $\mathcal{R}(\mathcal{M})=\mathcal{M}$. Figure 4(a) shows that kinetic theory successfully describes the optimization of the controller (theoretical calculations in SM [70]), by predicting the long-time behavior of $\mu(t)$. In the absence of mutations ($D_{M}=0$), the diversity $\sigma^{2}(t)$ goes to zero and $\mu(t)$ approaches the maximal reward around $\chi_{T}=1.09$ [see Fig. 4(b)]. However, for $D_{M}>0$, mutations cause the robots to learn a distribution of policies with $\sigma^{2}_{\infty}>0$. Importantly, the underlying model implies a pronounced asymmetry ¯$\bar{\mathcal{R}}(\chi_{T}+\epsilon)>\bar{\mathcal{R}}(\chi_{T}-\epsilon),\epsilon>0$ [Fig. 4(b)], where ¯$\bar{\mathcal{R}}(\chi)=\mathcal{R}(\mu_{\mathcal{M}}(\chi))$ is the average steady-state reward for an agent with policy $\chi$. This makes it preferable to adapt for $t\to\infty$ to $\mu_{\infty}>\chi_{T}$ in the presence of mutations, as observed in the agent-based simulations. Expanding $\mu_{\mathcal{M}}$ to fourth order in Eq. (7), our kinetic theory provides a quantitative prediction of $\mu_{\infty}$ [Fig. 4(a)].
Conclusion and outlook – We proposed a kinetic theory framework for decentralized learning in smart active matter, in which agents maximize their reward $\mathcal{R}$ by learning improved policies $\mathcal{P}$ from neighboring agents, to reach a target collective state. This approach enables the derivation of space-dependent field equations describing the time-dependence of policy distributions. This framework has been applied to two explicit models, and predictions favorably compared to numerical simulations of the agent-based model, already showing a rich phenomenology. Future research direction may include the analysis of learning with different interaction ranges, multi-dimensional policies and more complex reward functions. While we considered for simplicity a limit of separated time scales between physical, memory and learning dynamics, our framework allows for systematic improvements over this approximation, notably taking into account the effect of the communication range. In addition, it would also be of interest to investigate the generality of the uncertainty relation linking learning time and policy fluctuations.
Acknowledgements.
Acknowledgments – The authors are grateful to Olivier Dauchot for stimulating discussions. This work has been supported by MIAI@Grenoble Alpes, (ANR-19-P3IA-0003).